OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (-2, -3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

A. x - y = 1 B. x + y = 1 C. x - y = -5 D. x + y = 5

OpenStudy (anonymous):

slope of the required line is same as the line(parallel) given so m=1 use point slope form to get ur answer

OpenStudy (anonymous):

(-2, -3) y-(-3)=m(x-(-2)) m=1 solve it